OpenStudy (anonymous):

Solve the system using elimination how to find y? I found ex 3 but not y 3x-2y=1 2x+2y=14

OpenStudy (anonymous):

if \(x=3\) then since \(3x-2y=1\) you know \[3\times 3-2y=1\] or \[9-2y=1\] and you can solve that for \(y\)

OpenStudy (anonymous):

alternatively, solve \[2\times 3+2y=14\] or \[6+2y=14\] you will get the same answer

OpenStudy (anonymous):

I'm a little confused here could you explain from the beginning

OpenStudy (anonymous):

did you find \(x=3\) ?

OpenStudy (anonymous):

no how did you get that?

OpenStudy (anonymous):

i thought you said \(x=3\) up top

OpenStudy (anonymous):

oh yeah I found that

OpenStudy (anonymous):

ok then the simplest way i know how to say it is that if x is 3 then you can replace x by 3 in either of the two original equation

OpenStudy (anonymous):

that is why i wrote \[2\times 3+2y=14\] or \[2\times 3+2y=14\]

OpenStudy (anonymous):

I know the answer is 4 just not sure how i got it

OpenStudy (anonymous):

do you know how to solve \[6+2y=14\]?"

OpenStudy (anonymous):

you divide 6 by 6 and 6 by 2 or?

OpenStudy (anonymous):

subtract 6 then divide by 2

OpenStudy (anonymous):

okay I got it now thank you :)

OpenStudy (anonymous):

yw